Let f(x)=x^2-25 and g(x)=x+5, find a)(f-g)(x) and b)(f/g)(x)
Answer by fractalier(6550)

Okay from
f(x)=x^2-25 and g(x)=x+5
we have
(f-g)(x) = x^2 - 25 - (x + 5) = x^2 - x - 30
and
(f/g)(x) = (x^2 - 25) / (x + 5) = x - 5